OpenStudy (anonymous):

1)A shuttle bus to take you from your house to the airport charges a $20.00 flat fee plus $0.50 for each mile driven. If the cost of the ride is $30.00, how many miles did you travel? Equation: 20+0.50x= 30

OpenStudy (anonymous):

20 + 0.50x = 30 -20 -20 0.50x = 30 0.50x/0.50 = 30/0.50 x = 60 so 60 miles

OpenStudy (anonymous):

oops, made a mistake 0.50x = 10 x = 20 20 miles, not 60

OpenStudy (texaschic101):

Nice save...I was wondering if you were gonna catch that...lol

OpenStudy (anonymous):

it took me a second lol